OpenStudy (astrophysics):

What is the most interesting math problem you've encountered that you could share? Post away!

OpenStudy (amilapsn):

Do you know how to prove all triangles are isosceles?

imqwerty (imqwerty):

- the sum of two opposite sides and a diagonal of a quadrilateral is 20cm. The area of quadrilateral is 50cm^2. Find the length of the other diagonal. this question is 100% correct and no information is missing :) :D

OpenStudy (astrophysics):

Haha, that sounds surprisingly easy, but it's not..hmm

imqwerty (imqwerty):

XD

ganeshie8 (ganeshie8):

My all time favorite math problem is a proof: The quadratic reciprocity law. For distinct odd primes \(p,q\), show that \[\large (p/q)(q/p)=(-1)^{\frac{p-1}{2}\frac{q-1}{2}}\] (Gauss got so much obsessed with this problem and called it law ) https://en.wikipedia.org/wiki/Quadratic_reciprocity

hartnn (hartnn):

*

OpenStudy (thomas5267):

@imqwerty Does the quadrilateral has to be convex?

OpenStudy (thomas5267):

Would this be a valid quadrilateral under the scope of that question? 

imqwerty (imqwerty):

its not given in the question so u can take any quadrilateral.

OpenStudy (thomas5267):

"the sum of two opposite sides and a diagonal of a quadrilateral is 20cm." Which of the following is true? one side + opposite side + diagonal = 20 cm; or, one side + opposite side = diagonal = 20 cm

imqwerty (imqwerty):

1st one

OpenStudy (mathmath333):

recently reasoning questions are fascinating

OpenStudy (nincompoop):

cool!

OpenStudy (nincompoop):

does it work with other integer or Z+ values?

OpenStudy (nincompoop):

the boxes for astro

OpenStudy (nincompoop):

nevermind, I just checked it myself

ganeshie8 (ganeshie8):

@mukushla is the answer emptyset ?

ganeshie8 (ganeshie8):

the given equation can be rearranged as a^3 - b^3 = 199*200*ab (a-b)(a^2+ab+b^2)=199*200*ab (a-b)(a/b+b/a+1) = 199*200 since the right hand side is integer, it must be the case that a/b+b/a is also an integer only integer value of a/b+b/a is 2 and this doesn't satisfy the equation, so there are no solutions

OpenStudy (dan815):